OpenStudy (anonymous):

what is difference ? f(x1) - f(x0) ---------- x1 - x0 f(x1) - f(x0) lim ---------- x1->x0 x1 - x0

OpenStudy (ash2326):

In the first one \[ \text{x1 may or may not be equal to x0}\] but in the second one x1 is approaching x0 and it's actually a limit test

OpenStudy (anonymous):

are both "rate of change" ?

OpenStudy (ash2326):

No, first is not rate of change If you look carefully, the second one is actually f'(x0) which is slope of f at x0

OpenStudy (anonymous):

ash.... why first one is not rate of change..... whereas second one is rate of change...

OpenStudy (ash2326):

Okay just wait for 5 minutes. I'll reconfirm

OpenStudy (anonymous):

sure...........................

OpenStudy (ash2326):

Yeah Jatin both are rate of change

OpenStudy (ash2326):

But there is a difference between those two!!!

OpenStudy (anonymous):

then what is the difference.. i'm so confused... i know that the second one is differentiation.

OpenStudy (anonymous):

what is the difference ?

OpenStudy (ash2326):

If f is a linear function like f(x)= mx+c then both will give the same results, say we have y=f(x)=2x let x1= 3 and x0=1 so first one will give \[\frac{f((x1)-f(x0)}{x1-x0}=\frac{6-2}{3-1}=2\] now you find the second one it's f'(3)

OpenStudy (anonymous):

yah... if f is linear function then this is true .. but what is the difference i really need to know

OpenStudy (ash2326):

@jatinbansalhot the first one is the average rate of change the second one is instantaneous rate of change

OpenStudy (ash2326):

I'll explain with a diagram

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhh......................... omg.. that is what i wanted thanks thanks really thanks a lot :) :) :)

OpenStudy (ash2326):

Say we have f which is not a linear function

OpenStudy (anonymous):

sure if u explain with dia then really thanks :)

OpenStudy (ash2326):

Yeah. Now I'll mark two points (x1, f(x1)) and (x0, f(x0)) (x1, x0 are very close to each other) on this curve and the first one which is the average rate of change considers this function as a line and just give the slope of this line You can see the line 12 1 is (x0, f(x0)) 2 is (x1, f(x1)) The first relation gives the slope of this line or simply average rate of change